OpenStudy (anonymous):

Which is the better deal? A. 20 lb of potatoes for $64.80 B. 25 lb of potatoes for $80.50 C.30 lb of potatoes for $97.20 D. 45 lb of potatoes for $145.80

OpenStudy (jdoe0001):

divide the price by the pounds, see how much each lb is really costing you the better deal will be the cheaper price per lb

OpenStudy (jdoe0001):

don't forget to get butter too =)

OpenStudy (anonymous):

A. $64.80/20=$3.24 per pound B. $80.50/25=$3.22 per pound C. $97.20/30=$3.24 per pound D. $145.80/45=$3.24 per pound So B would be the best price at $3.22 per pound, compared to the other three which are 2 cents more expensive.

OpenStudy (anonymous):

thanks guys